Ticket Number: HEA-NOC/19981224-01 Ticket Status: CLOSE
Ticket Type: Unscheduled Ticket Source: HEA-NOC
Ticket Scope: Site Site/Line: ITnet
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start: 19981223 17:00 UTC
Ticket Closed: 19981225 15:15:32 UTC Problem End: 19981224 17:00 UTC
ITnet router at Tallaght IT stops advertising routes to HEAnet.
Router password changed after 15:30.
External (to ITnet) connectivity of all ITnet sites.
07:40 Report to ITnet support (LAN Comms). Kevin G will try to
contact Tallaght with a view to going on site.
17:00 Access gained to IT Tallaght; Lancomms intervened at the
router. Traffic is now being routed again. Continuing to monitor.
15:00 Dec 25th: All clear, closing ticket. Kevin G made final
adjustments on-site in Tallaght at 19:30 yesterday.
Time to Fix:
Approx 24 h. Additional Notes:
Thanks to Kevin Gannon of LAN Comms, Martin Gibbons of GMIT, and Paul
Campbell of IT Tallaght, who took time out over the holiday period
to see that the problem was solved.